The Auto Group has 1,000 bonds outstanding that are selling for $970 each. The company also has 9,600 shares of preferred stock at a market price of $80 each. The common stock Is priced at $75 a share and there are 42,000 shares outstanding. What Is the welght of the preferred stock as it relates to the firms weighted average cost of capital? 186 percen 19.84 percent 31.42 percent O 15.71 percent 64.44 percent
Answer #1

Market value of bonds = 1,000 * 970 = 970,000

Market value of preferred stock = 9,600 * 80 = 768,000

Market value of common stock = 42,000 * 75 = 3,150,000

total market value = 970,000 + 768,000 + 3,150,000 = 4,888,000

Weight of preferred stock = ( 768,000 / 4,888,000) * 100

Weight of preferred stock = 0.1571 or 15.71%